Use the discriminant to find how many places the quadratic function crosses the x-axis.
4x^2 + 8x = 0


a) Two zeros
b) One zero
c) No zeros

Answer by stanbon(75887)

Use the discriminant to find how many places the quadratic function crosses the x-axis.
4x^2 + 8x = 0
D = sqrt(8^2 - 4*4*0) = sqrt(8^2) = 8
Since D > 0 the function will have two zeros
=============
Cheers,
Stan H.

a) Two zeros
b) One zero
c) No zeros
0 solutions